# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 4.2. The principal $A$-determinant and nondegenerate polynomials
We do not need the following remark for what follows, but we mention the beautiful theorem [GKZ94, Theorem 10.1.4] that the Newton polytope of $E_{A}(z_{a})$ in $\mathbb{R}^{A}$ is naturally identified with the secondary polytope $\Sigma(A)$ of $A$ [DLRS10, Definition 5.1.6]. The faces $F(\mathcal{S},B)$ of the secondary polytope are in bijection with pairs $(\mathcal{S},B)$, where $\mathcal{S}$ is a regular (lattice) polyhedral subdivision of $P$ and $B\subset A$ contains the vertices of $\mathcal{S}$. The cone corresponding to $F(\mathcal{S},B)$ in the dual fan to $\Sigma(A)$ has relative interior $\{\omega:A\to\mathbb{R}:(\mathcal{S},B)=(\mathcal{S}(\omega),A_{\omega})\}$. The Newton polytope $\operatorname{Newt}(E_{A,M})$ is a dilate of $\operatorname{Newt}(E_{A})$ and hence is combinatorially equivalent. Then it follows from the definitions that $\operatorname{init}_{\omega}E_{A,M}$ is the restriction of $E_{A,M}$ to the face of $\operatorname{Newt}(E_{A,M})$ corresponding to $F(\mathcal{S}(\omega),A_{\omega})$. Finally, recall that we have an action $\rho:G\to\operatorname{Aff}(M)$ and $P$ is $G$-invariant. We will not need this in what follows, but we mention that when $A$ is $G$-invariant, Reiner introduced the equivariant secondary polytope of $A$ in [Rei02], with faces in bijection with all pairs $(\mathcal{S},B)$ above such that $\mathcal{S}$ and $B$ are $G$-invariant.
@@
$x^{u_{1}}\cdot x^{u_{2}}=\begin{cases}x^{u_{1}+u_{2}}&\textrm{ if there exists}C\in\Sigma\textrm{ such that }u_{1},u_{2}\in C,\\ 0&\textrm{ otherwise. }\end{cases}$
@@ 4.3. Applications to equivariant Ehrhart theory
Recall that $\widetilde{N}_{C}=\operatorname{Hom}_{\mathbb{C}}(\widetilde{M}_{\mathbb{C}},\mathbb{C})$ is naturally a $\mathbb{C}G$-module. Explicitly, the linear action $\rho:G\to\operatorname{GL}(\widetilde{M}_{\mathbb{C}})$ induces a linear action $\rho^{*}:G\to\operatorname{GL}(\widetilde{N}_{\mathbb{C}})$ determined by $\langle u,\rho^{*}(g)(v)\rangle=\langle\rho(g^{-1})(u),v\rangle$ for all $u$ in $\widetilde{M}_{\mathbb{C}}$ and $v$ in $\widetilde{N}_{\mathbb{C}}$. Consider the map of $\mathbb{C}$-vector spaces:
@@
We also claim that $\psi$ is injective. Suppose that $\partial_{v}F=0$ for some $v$ in $\widetilde{N}_{\mathbb{C}}$. Then $\langle(u,1),v\rangle=0$ for all $u\in\operatorname{supp}(f)$. Since $\operatorname{Newt}(f)=P$, the linear span of $\operatorname{supp}(f)\times\{1\}$ equals $\widetilde{M}_{\mathbb{C}}$ and hence $v=0$. We conclude that $\psi$ is an isomorphism of $\mathbb{C}G$-modules. It remains to observe that $\widetilde{N}_{\mathbb{C}}$ is isomorphic to $\widetilde{M}_{\mathbb{C}}$ as a $\mathbb{C}G$-module. This follows since if $\rho(g)$ has eigenvalues $\{\lambda_{1},\ldots,\lambda_{d+1}\}$, then $\rho^{*}(g)$ has eigenvalues $\{\lambda_{1}^{-1},\ldots,\lambda_{d+1}^{-1}\}$. On the other hand, since $\rho(g)$ is integer valued and each eigenvalue is a root of unity, $\lambda_{i}^{-1}$ is the complex conjugate of $\lambda_{i}$ and is also an eigenvalue of $\rho(g)$ for all $11\leq i\leq d+$ 1 (cf. the proof of Proposition 2.28). ∎
@@
By Theorem 4.18, $f|_{Q}$ is not nondegenerate with respect to $Q$. By Definition 4.25, $f$ is not nondegenerate with respect to $\mathcal{S}$. ∎
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $\widetilde{M}=M\oplus\mathbb{Z}$ and let $C_{P}$ be the cone over $P\times\{1\}$ in $\widetilde{M}_{\mathbb{R}}$ with corresponding graded semigroup algebra $S_{P}=\mathbb{C}[C_{P}\cap\widetilde{M}]$. Consider the following conditions:
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $[\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}]$ in $R(G)$.
@@
    There exists a $G$-invariant lattice polyhedral subdivision $\mathcal{S}$ of $P$ and a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{\mathcal{S}}$ such that $[\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}]$ in $R(G)$.
@@
Then the following implications hold: (1) normal-$\Rightarrow$ (2) normal-$\Rightarrow$ (3) normal-$\Rightarrow$ (4).
@@
In the examples presented in this paper, we will show that the answer to Question 4.33 is that all conditions of Corollary 4.32 are equivalent. The only exception is Example 2.16, where we leave this question open (cf. Question 3.21).
@@
Suppose there exists a $G$-invariant triangulation $\mathcal{S}$ of $P$ with vertices in $\frac{1}{N}M$. Recall that $S_{\mathcal{S}}=\oplus_{m\in\mathbb{Z}_{\geq 0}}S_{\mathcal{S},m}$ is the toric face ring associated to $\mathcal{S}$. Consider the subalgebra $S^{\prime}=\oplus_{m\in N\mathbb{Z}_{\geq 0}}S_{\mathcal{S},m}$ of $S_{\mathcal{S}}$. Then $S_{\mathcal{S}}$ is integral over $S^{\prime}$ and hence $S_{\mathcal{S}}$ is a finite $S^{\prime}$-module by Lemma 4.1. Recall that $\widetilde{M}=M\oplus\mathbb{Z}$ and $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ is projection onto the last coordinate, with corresponding linear map $\operatorname{pr}_{\mathbb{R}}:\widetilde{M}_{\mathbb{R}}\to\mathbb{R}$. As in Remark 2.12, we may consider the affine lattice $M^{\prime}=\operatorname{pr}^{-1}(N)$ in the real affine space $\operatorname{pr}_{\mathbb{R}}^{-1}(N)$, with corresponding affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$ and the $G$-invariant polytope $P^{\prime}:=C_{P}\cap\operatorname{pr}_{\mathbb{R}}^{-1}(N)$. In particular, the equivariant Hilbert series of $S^{\prime}$ is $\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$. The latter is naturally identified with $\operatorname{Ehr}(P,\rho_{N};t)$ where $\rho_{N}:G\to\operatorname{Aff}(\frac{1}{N}M)$ is the corresponding affine representation of $G$ on $\frac{1}{N}M$ considered in the introduction and Section 2.5.
As in the proof of Theorem 1.4, by Lemma 4.27 and Lemma 4.28 there exists a h.s.o.p. $F_{1},\ldots,F_{d+1}$ for $S^{\prime}$ of degree $N$ (here we view $S^{\prime}$ as graded by $N\mathbb{Z}$ rather than $\mathbb{Z}$) such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}$ is $G$-invariant and isomorphic to $\widetilde{M}_{\mathbb{C}}$ as a $\mathbb{C}G$-module. By Corollary 4.5, $F_{1},\ldots,F_{d+1}$ is a h.s.o.p. of degree $N$ of $S_{\mathcal{S}}$. By Lemma 4.8 and Theorem 4.24, $h^{*}_{N}(P,\rho;t^{N})=\operatorname{Hilb}_{G}(S_{\mathcal{S}}/(F_{1},\ldots,F_{d+1}))$. In particular, $h^{*}_{N}(P,\rho;t)$ is effective. Finally, if $h^{*}_{N}(P,\rho;t)$ is effective, then $h^{*}_{N}(P,\rho;t^{N})(1)=h^{*}(P;t)(1+t+\cdots+t^{N-1})^{d+1}$ has nonnegative integer coefficients and encodes the dimensions of the representations of the coefficients of $h^{*}_{N}(P,\rho;t^{N})$. In particular, $h^{*}_{N}(P,\rho;t)$ is a polynomial (cf. Remark 2.13). ∎
@@
Let $A$ be the set of vertices of $\mathcal{S}$, and let $f=\sum_{a\in A}x^{a}\in\mathbb{C}[M]$. If $F=\sum_{a\in A}x^{(a,1)}$ is the element in $S_{\mathcal{S},1}$ corresponding to $f$, then $p(F)$ is the element in $(S_{\mathcal{S}|_{Q}})_{1}$ corresponding to $f|_{Q}$. Let $\widetilde{M}_{Q}$ be the intersection of $\widetilde{M}$ with the linear span of $C_{Q}$. For any $v\in\widetilde{N}_{\mathbb{C}}=\operatorname{Hom}_{\mathbb{C}}(\widetilde{M}_{\mathbb{C}},\mathbb{C})$, we have $p(\partial_{v}F)=\partial_{v_{Q}}p(F)$, where $v_{Q}$ is the restriction of $v$ to $(\widetilde{M}_{Q})_{\mathbb{C}}$. It follows that $p(J_{f,\mathcal{S}})=J_{f|_{Q},\mathcal{S}|_{Q}}$, and $p$ induces a map of graded $\mathbb{C}G$-modules $\bar{p}:S_{\mathcal{S}}/J_{f,\mathcal{S}}\to S_{\mathcal{S}|_{Q}}/J_{f|_{Q},\mathcal{S}|_{Q}}.$ By Lemma 4.28, $h^{*}(P,\rho;t)=\operatorname{Hilb}_{G}(S_{\mathcal{S}}/J_{f,\mathcal{S}};t)$ and $h^{*}(P,\rho_{Q};t)=\operatorname{Hilb}_{G}(S_{\mathcal{S}|_{Q}}/J_{f|_{Q},\mathcal{S}|_{Q}};t)$. The result now follows from the fact that $\bar{p}$ is surjective.
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $[\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}]$ in $R(G)$.
@@
Theorem 1.4 gives the implications (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5). Let $A=\{x_{1},\ldots,x_{d+2}\}$. Configurations of $d+2$ points that affinely span a $d$-dimensional space are well studied, and we refer the reader to [DLRS10, Section 2.4.1] and [GKZ94, Section 7.1.B] for details. There is a unique (up to scaling) affine relation between the elements of $A$. This implies that there exists unique positive integers $a_{1},\ldots,a_{\ell}$ and $a_{1}^{\prime},\ldots,a_{\ell^{\prime}}^{\prime}$ such that $\ell+\ell^{\prime}\leq d+2$, $\sum_{i=1}^{\ell}a_{i}=\sum_{i=1}^{\ell^{\prime}}a_{i}^{\prime}$, $\gcd(a_{1},\ldots,a_{\ell},a_{1}^{\prime},\ldots,a_{\ell^{\prime}}^{\prime})=1$ and $a_{1}x_{1}+\cdots+a_{\ell}x_{\ell}-(a_{1}^{\prime}x_{\ell+1}+\cdots+a_{\ell^{\prime}}^{\prime}x_{d+2})=0$. By [DLRS10, Lemma 2.4.2], there exists precisely two (regular) lattice triangulations of $P$ with vertices in $A$: $T$ with facets $\{\operatorname{Conv}\left\{{A\smallsetminus\{x_{i}\}}\right\}:1\leq i\leq\ell\}$, and $T^{\prime}$ with facets $\{\operatorname{Conv}\left\{{A\smallsetminus\{x_{i}\}}\right\}:\ell+1\leq i\leq\ell+\ell^{\prime}\}$. Since $G$ acts permuting $A$, $G$ acts on $\{T,T^{\prime}\}$. By assumption, there does not exist a $G$-invariant triangulation of $P$, and hence there exists $g$ in $G$ such that $g\cdot T=T^{\prime}$. Hence $\ell=\ell^{\prime}$ and, after possibly reindexing coordinates, we may assume that $g\cdot x_{i}=x_{\ell+i}$ and $a_{i}=a_{i}^{\prime}$ for $1\leq i\leq\ell$. By [GKZ94, Theorem 10.1.4] and [GKZ94, Theorem 10.1.14], the Newton polytope of $E_{A}$ (equal to the secondary polytope of $A$ by Remark 4.22) is isomorphic to a unit interval, and up to multiplication by a nonzero scalar and a monomial, $E_{A}(z_{1},\ldots,z_{d+1})$ equals $\prod_{i=1}^{\ell}z_{i}-(-1)^{\operatorname{Vol}(P)}\prod_{i=1}^{\ell}z_{\ell+i}$. In particular, if $\operatorname{Vol}(P)$ is odd, then $E_{A}(1,\ldots,1)\neq 0$. By Theorem 4.18, we deduce that (1) $\Rightarrow$ (2). It remains to show that (5) $\Rightarrow$ (1). Note that $Q=\operatorname{Conv}\left\{{x_{1},\ldots,x_{2\ell}}\right\}$ is the unique minimal face $Q$ of $P$ that is not a simplex. Note that $Q$ is $G$-invariant. Since $A$ affinely generates $M$, it follows that $P$ is isomorphic to a free join of $Q$ and a standard simplex. In particular, $\operatorname{Vol}(P)=\operatorname{Vol}(Q)$ and Lemma 2.22 implies that the equivariant $h^{*}$-polynomials of $P$ and $Q$ agree. After possibly replacing $P$ with $Q$, we reduce to the case when $22\ell=d+$ 2. Since $A$ affinely generates $M$, then by regarding $M$ as an affine lattice we may assume that we are in the case of Example 2.17. Consider an element $g$ in $G$ such that $g\cdot T=T^{\prime}$. By Example 2.17, $h^{*}(P,\rho;t)(g)$ is a polynomial if and only if $\operatorname{Vol}(P)$ is odd, and we deduce that (5) $\Rightarrow$ (1).
@@
With the notation of Example 2.17, let $d=4$, $r=3$ and $a_{1}=a_{2}=a_{3}=1$. That is, let Let $e_{1},e_{2},e_{3},f_{1},f_{2},f_{3}$ be a basis of $\mathbb{Z}^{6}$, and let $\bar{e}_{i}$ and $\bar{f_{i}}$ denote the images of $e_{i}$ and $f_{i}$ respectively in the lattice $L:=\mathbb{Z}^{6}/\mathbb{Z}(e_{1}+e_{2}+e_{3}-f_{1}-f_{2}-f_{3})$. Consider the group homomorphism $\psi:L\to\mathbb{Z}$ satisfying $\psi(\bar{e}_{i})=\psi(\bar{f}_{i})=1$ for all $1\leq i\leq r$. As in Example 2.4, $M^{\prime}=\psi^{-1}(1)$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Let $P=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{2},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2},\bar{f}_{3}}\right\}\subset W$. Then $P$ is a $4$-dimensional lattice polytope with $6$ vertices and $h^{*}(P;t)=1+t+t^{2}$. Let $G=\mathbb{Z}/2\mathbb{Z}=\langle\sigma\rangle$ and consider the affine representation $\rho:G\to\operatorname{Aff}(M^{\prime})$ determined by $\sigma\cdot\bar{e}_{i}=\bar{f}_{i}$ for $1\leq i\leq 3$. Since the only lattice points of $P$ are its $d+2$ vertices, as in the proof of Proposition 4.34, $P$ has precisely two (regular) lattice triangulations $T,T^{\prime}$, and $\sigma\cdot T=T^{\prime}$. In particular, there does not exist a $G$-invariant lattice triangulation of $P$. Since $\operatorname{Vol}(P)=3$ is odd, Proposition 4.34 implies that there exists a $G$-invariant polynomial that is nondegenerate with respect to $P$. Moreover, by Example 2.17, $h^{*}(P,\rho;t)(\sigma)=1-t+t^{2}$, and hence $h^{*}(P,\rho;t)=1+\chi t+t^{2}$, where $\chi$ is the nontrivial representation of $G$ i.e. $\chi(\sigma)=-1$. In particular, since the multiplicity of the trivial representation in the (nonzero) linear term of $h^{*}(P,\rho;t)$ is zero, this gives a counterexample to Conjecture 1.3.

### v3

@@ 4.2. The principal $A$-determinant and nondegenerate polynomials
We do not need the following remark for what follows, but we mention the beautiful theorem [GKZ94, Theorem 10.1.4] that the Newton polytope of $E_{A}(z_{a})$ in $\mathbb{R}^{A}$ is naturally identified with the secondary polytope $\Sigma(A)$ of $A$ [DLRS10, Definition 5.1.6]. The faces $F(\mathcal{S},B)$ of the secondary polytope are in bijection with pairs $(\mathcal{S},B)$, where $\mathcal{S}$ is a regular (lattice) polyhedral subdivision of $P$ and $B\subset A$ contains the vertices of $\mathcal{S}$. The cone corresponding to $F(\mathcal{S},B)$ in the dual fan to $\Sigma(A)$ has relative interior $\{\omega:A\to\mathbb{R}:(\mathcal{S},B)=(\mathcal{S}(\omega),A_{\omega})\}$.
The Newton polytope $\operatorname{Newt}(E_{A,M})$ is a dilate of $\operatorname{Newt}(E_{A})$ and hence is combinatorially equivalent. Then it follows from the definitions that $\operatorname{init}_{\omega}E_{A,M}$ is the restriction of $E_{A,M}$ to the face of $\operatorname{Newt}(E_{A,M})$ corresponding to $F(\mathcal{S}(\omega),A_{\omega})$.
Finally, recall that we have an action $\rho:G\to\operatorname{Aff}(M)$ and $P$ is $G$-invariant. We will not need this in what follows, but we mention that when $A$ is $G$-invariant, Reiner introduced the equivariant secondary polytope of $A$ in [Rei02], with faces in bijection with all pairs $(\mathcal{S},B)$ above such that $\mathcal{S}$ and $B$ are $G$-invariant.
@@
[…]
@@ 4.3. Applications to equivariant Ehrhart theory
[…]
@@
We also claim that $\psi$ is injective. Suppose that $\partial_{v}F=0$ for some $v$ in $\widetilde{N}_{\mathbb{C}}$. Then $\langle(u,1),v\rangle=0$ for all $u\in\operatorname{supp}(f)$. Since $\operatorname{Newt}(f)=P$, the linear span of $\operatorname{supp}(f)\times\{1\}$ equals $\widetilde{M}_{\mathbb{C}}$ and hence $v=0$. We conclude that $\psi$ is an isomorphism of $\mathbb{C}G$-modules.
It remains to observe that $\widetilde{N}_{\mathbb{C}}$ is isomorphic to $\widetilde{M}_{\mathbb{C}}$ as a $\mathbb{C}G$-module. This follows since if $\rho(g)$ has eigenvalues $\{\lambda_{1},\ldots,\lambda_{d+1}\}$, then $\rho^{*}(g)$ has eigenvalues $\{\lambda_{1}^{-1},\ldots,\lambda_{d+1}^{-1}\}$. On the other hand, since $\rho(g)$ is integer valued and each eigenvalue is a root of unity, $\lambda_{i}^{-1}$ is the complex conjugate of $\lambda_{i}$ and is also an eigenvalue of $\rho(g)$ for all $1\leq i\leq d+1$ (cf. the proof of Proposition 2.28). ∎
@@
By Theorem 4.18, $f|_{Q}$ is not nondegenerate with respect to $Q$. By Definition 4.25, $f$ is not nondegenerate with respect to $\mathcal{S}$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $\widetilde{M}=M\oplus\mathbb{Z}$ and let $C_{P}$ be the cone over $P\times\{1\}$ in $\widetilde{M}_{\mathbb{R}}$ with corresponding graded semigroup algebra $S_{P}=\mathbb{C}[C_{P}\cap\widetilde{M}$]. Consider the following conditions:
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}$] in $R(G)$.
@@
    There exists a $G$-invariant lattice polyhedral subdivision $\mathcal{S}$ of $P$ and a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{\mathcal{S}}$ such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}$] in $R(G)$.
@@
Then the following implications hold: (1) $\Rightarrow$ (2) $\Rightarrow$ (3) $\Rightarrow$ (4).
@@
In the examples presented in this paper, we will show that the answer to Question 4.33 is that all conditions of Corollary 4.32 are equivalent. The only exception is Example 2.16, where we leave this question open (cf. Question 3.22).
@@
Suppose there exists a $G$-invariant triangulation $\mathcal{S}$ of $P$ with vertices in $\frac{1}{N}M$. Recall that $S_{\mathcal{S}}=\oplus_{m\in\mathbb{Z}_{\geq 0}}S_{\mathcal{S},m}$ is the toric face ring associated to $\mathcal{S}$. Consider the subalgebra $S^{\prime}=\oplus_{m\in N\mathbb{Z}_{\geq 0}}S_{\mathcal{S},m}$ of $S_{\mathcal{S}}$. Then $S_{\mathcal{S}}$ is integral over $S^{\prime}$ and hence $S_{\mathcal{S}}$ is a finite $S^{\prime}$-module by Lemma 4.1.
Recall that $\widetilde{M}=M\oplus\mathbb{Z}$ and $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ is projection onto the last coordinate, with corresponding linear map $\operatorname{pr}_{\mathbb{R}}:\widetilde{M}_{\mathbb{R}}\to\mathbb{R}$. As in Remark 2.12, we may consider the affine lattice $M^{\prime}=\operatorname{pr}^{-1}(N)$ in the real affine space $\operatorname{pr}_{\mathbb{R}}^{-1}(N)$, with corresponding affine representation $\rho^{\prime}:G\to\operatorname{Aff}(M^{\prime})$ and the $G$-invariant polytope $P^{\prime}:=C_{P}\cap\operatorname{pr}_{\mathbb{R}}^{-1}(N)$. In particular, the equivariant Hilbert series of $S^{\prime}$ is $\operatorname{Ehr}(P^{\prime},\rho^{\prime};t)$. The latter is naturally identified with $\operatorname{Ehr}(P,\rho_{N};t)$ where $\rho_{N}:G\to\operatorname{Aff}(\frac{1}{N}M)$ is the corresponding affine representation of $G$ on $\frac{1}{N}M$ considered in the introduction and Section 2.5.
As in the proof of Theorem 1.4, by Lemma 4.27 and Lemma 4.28 there exists a h.s.o.p. $F_{1},\ldots,F_{d+1}$ for $S^{\prime}$ of degree $N$ (here we view $S^{\prime}$ as graded by $N\mathbb{Z}$ rather than $\mathbb{Z}$) such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}$ is $G$-invariant and isomorphic to $\widetilde{M}_{\mathbb{C}}$ as a $\mathbb{C}G$-module. By Corollary 4.5, $F_{1},\ldots,F_{d+1}$ is a h.s.o.p. of degree $N$ of $S_{\mathcal{S}}$. By Lemma 4.8 and Theorem 4.24, $h^{*}_{N}(P,\rho;t^{N})=\operatorname{Hilb}_{G}(S_{\mathcal{S}}/(F_{1},\ldots,F_{d+1}))$. In particular, $h^{*}_{N}(P,\rho;t)$ is effective.
Finally, if $h^{*}_{N}(P,\rho;t)$ is effective, then $h^{*}_{N}(P,\rho;t^{N})(1)=h^{*}(P;t)(1+t+\cdots+t^{N-1})^{d+1}$ has nonnegative integer coefficients and encodes the dimensions of the representations of the coefficients of $h^{*}_{N}(P,\rho;t^{N})$. In particular, $h^{*}_{N}(P,\rho;t)$ is a polynomial (cf. Remark 2.13).
@@
[…]
@@
    There exists a l.s.o.p. $F_{1},\ldots,F_{d+1}$ of $S_{P}$ such that $\mathbb{C}F_{1}+\cdots+\mathbb{C}F_{d+1}]=[\widetilde{M}_{\mathbb{C}}$] in $R(G)$.
@@
Theorem 1.4 gives the implications (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5). Let $A=\{x_{1},\ldots,x_{d+2}\}$. Configurations of $d+2$ points that affinely span a $d$-dimensional space are well studied, and we refer the reader to [DLRS10, Section 2.4.1] and [GKZ94, Section 7.1.B] for details. There is a unique (up to scaling) affine relation between the elements of $A$. This implies that there exists unique positive integers $a_{1},\ldots,a_{\ell}$ and $a_{1}^{\prime},\ldots,a_{\ell^{\prime}}^{\prime}$ such that $\ell+\ell^{\prime}\leq d+2$, $\sum_{i=1}^{\ell}a_{i}=\sum_{i=1}^{\ell^{\prime}}a_{i}^{\prime}$, $\gcd(a_{1},\ldots,a_{\ell},a_{1}^{\prime},\ldots,a_{\ell^{\prime}}^{\prime})=1$ and $a_{1}x_{1}+\cdots+a_{\ell}x_{\ell}-(a_{1}^{\prime}x_{\ell+1}+\cdots+a_{\ell^{\prime}}^{\prime}x_{d+2})=0$. By [DLRS10, Lemma 2.4.2], there exists precisely two (regular) lattice triangulations of $P$ with vertices in $A$: $T$ with facets $\{\operatorname{Conv}\left\{{A\smallsetminus\{x_{i}\}}\right\}:1\leq i\leq\ell\}$, and $T^{\prime}$ with facets $\{\operatorname{Conv}\left\{{A\smallsetminus\{x_{i}\}}\right\}:\ell+1\leq i\leq\ell+\ell^{\prime}\}$. Since $G$ acts permuting $A$, $G$ acts on $\{T,T^{\prime}\}$. By assumption, there does not exist a $G$-invariant triangulation of $P$, and hence there exists $g$ in $G$ such that $g\cdot T=T^{\prime}$. Hence $\ell=\ell^{\prime}$ and, after possibly reindexing coordinates, we may assume that $g\cdot x_{i}=x_{\ell+i}$ and $a_{i}=a_{i}^{\prime}$ for $1\leq i\leq\ell$.
By [GKZ94, Theorem 10.1.4] and [GKZ94, Theorem 10.1.14], the Newton polytope of $E_{A}$ (equal to the secondary polytope of $A$ by Remark 4.22) is isomorphic to a unit interval, and up to multiplication by a nonzero scalar and a monomial, $E_{A}(z_{1},\ldots,z_{d+1})$ equals $\prod_{i=1}^{\ell}z_{i}-(-1)^{\operatorname{Vol}(P)}\prod_{i=1}^{\ell}z_{\ell+i}$. In particular, if $\operatorname{Vol}(P)$ is odd, then $E_{A}(1,\ldots,1)\neq 0$. By Theorem 4.18, we deduce that (1) $\Rightarrow$ (2).
It remains to show that (5) $\Rightarrow$ (1). Note that $Q=\operatorname{Conv}\left\{{x_{1},\ldots,x_{2\ell}}\right\}$ is the unique minimal face $Q$ of $P$ that is not a simplex. Note that $Q$ is $G$-invariant. Since $A$ affinely generates $M$, it follows that $P$ is isomorphic to a free join of $Q$ and a standard simplex. In particular, $\operatorname{Vol}(P)=\operatorname{Vol}(Q)$ and Lemma 2.22 implies that the equivariant $h^{*}$-polynomials of $P$ and $Q$ agree. After possibly replacing $P$ with $Q$, we reduce to the case when $2\ell=d+2$. Since $A$ affinely generates $M$, then by regarding $M$ as an affine lattice we may assume that we are in the case of Example 2.17. Consider an element $g$ in $G$ such that $g\cdot T=T^{\prime}$. By Example 2.17, $h^{*}(P,\rho;t)(g)$ is a polynomial if and only if $\operatorname{Vol}(P)$ is odd, and we deduce that (5) $\Rightarrow$ (1).
@@
With the notation of Example 2.17, let $d=4$, $r=3$ and $a_{1}=a_{2}=a_{3}=1$. That is, let Let $e_{1},e_{2},e_{3},f_{1},f_{2},f_{3}$ be a basis of $\mathbb{Z}^{6}$, and let $\bar{e}_{i}$ and $\bar{f_{i}}$ denote the images of $e_{i}$ and $f_{i}$ respectively in the lattice $L:=\mathbb{Z}^{6}/\mathbb{Z}(e_{1}+e_{2}+e_{3}-f_{1}-f_{2}-f_{3})$. Consider the group homomorphism $\psi:L\to\mathbb{Z}$ satisfying $\psi(\bar{e}_{i})=\psi(\bar{f}_{i})=1$ for all $1\leq i\leq r$. As in Example 2.4, $M^{\prime}=\psi^{-1}(1)$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Let $P=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{2},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2},\bar{f}_{3}}\right\}\subset W$. Then $P$ is a $4$-dimensional lattice polytope with $6$ vertices and $h^{*}(P;t)=1+t+t^{2}$. Let $G=\mathbb{Z}/2\mathbb{Z}=\langle\sigma\rangle$ and consider the affine representation $\rho:G\to\operatorname{Aff}(M^{\prime})$ determined by $\sigma\cdot\bar{e}_{i}=\bar{f}_{i}$ for $1\leq i\leq 3$. Since the only lattice points of $P$ are its $d+2$ vertices, as in the proof of Proposition 4.34, $P$ has precisely two (regular) lattice triangulations $T,T^{\prime}$, and $\sigma\cdot T=T^{\prime}$. In particular, there does not exist a $G$-invariant lattice triangulation of $P$. Since $\operatorname{Vol}(P)=3$ is odd, Proposition 4.34 implies that there exists a $G$-invariant polynomial that is nondegenerate with respect to $P$.
Moreover, by Example 2.17, $h^{*}(P,\rho;t)(\sigma)=1-t+t^{2}$, and hence $h^{*}(P,\rho;t)=1+\chi t+t^{2}$, where $\chi$ is the nontrivial representation of $G$ i.e. $\chi(\sigma)=-1$. In particular, since the multiplicity of the trivial representation in the (nonzero) linear term of $h^{*}(P,\rho;t)$ is zero, this gives a counterexample to Conjecture 1.3.